OpenStudy (anonymous):

Find, correct to 1 decimal place, the two smallest positive values of θ which satisfy each of the following equations. (a) sin θ° = 0.1 (b) sin θ° = -0.84 (c) sin θ° = -0.951

OpenStudy (mathmale):

The first equation is easiest to work with, because the sine is positive. Ask yourself: "In which quadrants is the sine positive? How is the sine definited? It's the ratio of the _____ to the ________ .

OpenStudy (anonymous):

first and second

OpenStudy (anonymous):

I have zero idea on how to solve these sorts of problems. All I know is the unit circle, x is cosine, y is sine, and tangent is cosine over sine.

OpenStudy (anonymous):

I mean sine over cosine

OpenStudy (mathmale):

How do you find the material to study this topic? Do you have a textbook? Are concepts explained on some web site to which you have access? Definitions of the sine, cosine and tangent functions are of critical importance in trig. I will use Equation Editor (below) to define them symbolically. Note that "opp" means opposite side, and that I will draw a triangle to illustrate this.

OpenStudy (mathmale):

"adj" means "adjacent side." "hyp" stands for "hypotenuse".

OpenStudy (mathmale):

\[\sin \theta = \frac{ opp }{ hyp}; \cos \theta = \frac{ adj }{ hyp }; \tan \theta=\frac{ opp }{ adj}\]

OpenStudy (mathmale):

See these definitions before?